Choose an expression that represents the perimeter of the rectangle shown.
Length=x-2
Width=x-4

Answer by jim_thompson5910(35256)

P = 2*(L+W)

P = 2*(x-2+x-4)

P = 2*(2x-6)

P = 2*(2x)+2(-6)

P = 4x - 12